The indentfirst package^{∗}

David Carlisle carlisle@cs.man.ac.uk

1995/11/23

Abstract

Make the first line of all sections etc., be indented by the usual paragraph indentation. This should work with all the standard document classes.

∗This file has version number v1.03, last revised 1995/11/23.